V= 30.0ml Concentration= 0.1001 M

1.) Using the volume and concentration of Ca(OH)2 recorded in your experiment, calculate the

volume of 0.0750 M phosphoric acid required to exactly react with it. This is the equivalence

point of your titration.

2.) Calculate the percent error of volumes at the equivalence point
Answer #1

equation of reaction :

3 Ca(OH)2 + 2 H3PO4 Ca3(PO4)2 + 6 H2O

Moles of Ca(OH)2 = Molairity * Volume = 0.1001 * 30 * 10-3 = 3.003 * 10-3 moles

Moles of H3PO4 = Moles of Ca(OH)2 * 2/3 (from the equation) = Molarity * Volume

3.003 * 10-3 * 2 / 3 = 0.0750 * V

Solving, we get: V = 26.693 ml